OpenStudy (anonymous):

Solve this system of equations using the addition method. x - y = 6 -3y = x + 22 @Math2400 @MAYSTEPH @mathstudent55

OpenStudy (anonymous):

x-y=6 -x-3y=22 --------- -4y=28 y=28/-4 y=-7 x - y = 6 x-7=6 x=6+7 x=13

OpenStudy (anonymous):

Type the correct answer in ordered pair form

OpenStudy (whpalmer4):

No, @shorty17, you've got the answer in front of you, you figure out how to type the correct answer in ordered pair form. Do you know what it is?

OpenStudy (anonymous):

no but it doenst get it

OpenStudy (whpalmer4):

an ordered pair looks like (x,y)

OpenStudy (anonymous):

Do you have any options?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

LOL okay :P

OpenStudy (anonymous):

k